OpenStudy (dtan5457):

Given a2=-18, a5=2/3, get a6

OpenStudy (dtan5457):

completely blanked out and forgot how to do this

OpenStudy (anonymous):

\[a_2=-18\\ a_3=a^2r\\ a^4=a_2r^2\\ a_5=a_2r^3\] put \[-18\times r^3=\frac{2}{3}\] and solve for \(r\)

OpenStudy (anonymous):

typo there, i meant \(a_3=a_2r\)

OpenStudy (dtan5457):

no use of any formulas can be used here?

OpenStudy (dtan5457):

to get a1 or r?